OpenStudy (calculusxy):

Medal!!! A 1580 kg car is traveling with a speed of 15.0 m / s. What is the net force that is required to bring the car to a halt in a distance of 50.0m?

OpenStudy (calculusxy):

@hartnn

OpenStudy (calculusxy):

Sorry it is not a closed question.

OpenStudy (welshfella):

use the work/energy principle work done in stopping the car = energy lost by the car

OpenStudy (calculusxy):

We're given a formula that has \[F = mass \times acceleration\] So we are supposed to use Newton's formula and figure the answer out.

hartnn (hartnn):

so you'll need the deceleration required to bring the car to the halt then you can just multiply it with the mass 1580

OpenStudy (welshfella):

yes - you can do it that way too

OpenStudy (calculusxy):

How would Figure out the deceleration? I am only given 15 m / s ?

hartnn (hartnn):

try this : \(v^2 = u^2 +2as \) v = final velocity = 0 u = initial velocity = 15 s = displacement = 50

OpenStudy (welshfella):

deceleration = rate of change of velocity

OpenStudy (calculusxy):

Ok give me a moment :)

OpenStudy (calculusxy):

So what is "a" in the formula? The acceleration?

hartnn (hartnn):

yep, the acceleration and if it comes out to be negative, then deceleration .

OpenStudy (calculusxy):

Okay

OpenStudy (calculusxy):

I am a bit confused now: So I have the equation w/ substitution: \[0 = 15^2 + 2a50\] The part where it has 2a50 doesn't make sense. So is it like 2(a + 50)?

hartnn (hartnn):

\(2\times a\times s \\ 2 \times a \times 50 \\ 100\times a = 100a\)

OpenStudy (calculusxy):

Oh okay.

OpenStudy (calculusxy):

\[0 = 15^2 + 100a\] \[0 = 225 + 100a\]\[-225 = 100a \] \[a = -2.25?\]

hartnn (hartnn):

thats right :) so just multiply it with mass to get the force required..

OpenStudy (calculusxy):

Wow!!! This was so much easier! Okay, so it will be: \[-2.25 \times 1580 = 3555\]

OpenStudy (calculusxy):

Am I correct?

hartnn (hartnn):

\(\huge \color{green} \checkmark \)

OpenStudy (calculusxy):

Thank you!!!!! You're such a great person !! :) Can you help me with like two more??

OpenStudy (calculusxy):

Yes or no???

hartnn (hartnn):

You're a great learner! and sure :)

OpenStudy (calculusxy):

Thanks I am creating a new post.